OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 2x + 1 >= 19? x >= 9 9 <= x A number line with a closed circle on 9 and shading to the right. A number line with a closed circle on 9 and shading to the left.

OpenStudy (anonymous):

\[2x+1\geq 19 \implies x\geq \frac{19-1}{2}=2\] That means that \[x \geq 9\] is false.

OpenStudy (anonymous):

*sigh* \[\frac{18}{2}\neq 2\] I bet you can calculate it yourself.